OpenStudy (anonymous):

If vector u = (5, 3) and vector v = (-1, 4), what is the component form of vector u + v?

OpenStudy (anonymous):

All you need to do for this type of problem is add the individual components. u+v = (5+(-1),3+4) = (4,7)